OpenStudy (anonymous):

Find the equation of the line with x − intercept (4,0) that is perpendicular to the line 3x − 8 y = 5 .

OpenStudy (anonymous):

set one of the variables equal to zero then find the other variable and repeat for x and y, then plug it into the equation for a line y=mx+b

OpenStudy (anonymous):

is this correct? please help me :)

OpenStudy (anonymous):

Hey @Brent0423 - so, to find perpendicular lines, you need to remember that the slope of the line is the *inverse* of the original line. So, you should put your original equation into slope intercept format (y=mx+b), and see if you can figure out the next step from there.

OpenStudy (anonymous):

Im confused which original equation?

OpenStudy (anonymous):

Your 3x-8y=5 - You need to put that into the y=mx+b format

OpenStudy (anonymous):

Ok so in y=mx+b Y stands for the y-coord M stands for slope X stands for my x-coord B stands for what???

OpenStudy (anonymous):

B stands for the 'y-intercept' Basically, this is where the line crosses the y-axis (that's the vertical one)

OpenStudy (anonymous):

Please show me what it would be in y=mx+b form

OpenStudy (anonymous):

No problem, so, let's look at your original equation: 3x-8y=5 We want to put Y on its own side of the equation, so subtract 3X from both sides, this gives you: -8y = -3x + 5 Now, I need to divide by (-8) in order to get y on its own, giving you: \[y=3/8 x + -5/8\]

OpenStudy (anonymous):

I got it!! Now what??

OpenStudy (anonymous):

So, what's your slope?

OpenStudy (anonymous):

3/8. So the negative reciprocal is -8/3

OpenStudy (anonymous):

^ Awesome!

OpenStudy (anonymous):

K, so now we need to give the full equation of the line. You've already got the slope, now can you figure out how to get b? Use the points they give you and put them all into the slope intercept equation...

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (anonymous):

y-0=-8/3(x-4) y=-8/3x+32/3

OpenStudy (anonymous):

I do believe that's the correct answer - but I'd like to make sure with @Hero first. Good work though @Brent0423 - way to stick in there :)